please help me solve 23/40 is what percent
Answer by stanbon(75887)

23/40 is what percent
-------
x/100 = 23/40
----
x = 100(23/40)
----
x = (5/2)23
----
x = 115/2
-----
x = 57.5%
============
Cheers,
Stan H.
============